a wire is need to support a vertical pole 14 feet high. the cable will be anchored to a stake 10 feet from the base of the pole. how much cable is needed?
Answer by mananth(16949)

cable , pole and the ground form a right triangle.
cable = hypotenuse
..
14^2+10^2=hyp^2
196+100=hyp^2
296=hyp^2
hyp= sqrt 296
17.2 m of cable